OpenStudy (geekfromthefutur):

PLZ help me ive tried every solution and i cant figure it out I WILL GIVE MEDAL AND FAN Evaluate. cube root of 7 multiplied by the square root of 7 over sixth root of 7 to the power of 5

OpenStudy (anonymous):

\[\frac{ \sqrt[3]{7}\times \sqrt[2]{7} }{ \sqrt[6]{7^{5}} }\] Is this what it looks like?